perimeter is 38'
sides: 15 -2x
length x +7
What are the dimensions

Answer by checkley71(8403)

P=2W+2L OR 38=2(15-2X)+2(X+7) OR 38=30-4X+2X+14 OR 38=44-2X OR 2X=6 OR X=3 THUS
W=15-2*3 OR W=15-6 OR W=9 & L=3+7 OR L=10
PROOF 38=2*9+2*10 OR 38=18+20 OR 38=38